# ClaimGen-CN: A Large-scale Chinese Dataset for Legal Claim Generation

### v1

@@ Abstract
Legal claims refer to the plaintiff’s demands in a case and are essential to guiding judicial reasoning and case resolution. While many works have focused on improving the efficiency of legal professionals, the research on helping non-professionals (e.g., plaintiffs) remains unexplored. This paper explores the problem of legal claim generation based on the given case’s facts. First, we construct ClaimGen-CN, the first dataset for Chinese legal claim generation task, from various real-world legal disputes. Additionally, we design an evaluation metric tailored for assessing the generated claims, which encompasses two essential dimensions: factuality and clarity. Building on this, we conduct a comprehensive zero-shot evaluation of state-of-the-art general and legal-domain large language models. Our findings highlight the limitations of the current models in factual precision and expressive clarity, pointing to the need for more targeted development in this domain. To encourage further exploration of this important task, we will make the dataset publicly available.
@@ 1 Introduction
  • To motivate other scholars to investigate this novel and important problem, we make the ClaimGen-CN dataset publicly available [^1] https://github.com/JosieZhou00/ClaimGen-CN.
@@ 7 Conclusion
The corpus we use is released by the Chinese government and has been anonymized wherever necessary [^5] Provisions of the Supreme People’s Court on Publication of Judgment Documents by the People’s Courts on the Internet 2016. http://gongbao.court.gov.cn/Details/415f49dd8baaa04b479d57af9616ef.html. Our dataset thus does not involve any personally private information. Besides, the corpus is in public domain and licensed for use within the legal scope [^6] See Article 3.3 of User Protocol of China Judgments Online. https://g.alicdn.com/onlineCourt/static/0.6.97/akan-wenshu-protocol.html.
@@
Measures for releasing the dataset: Our dataset is sourced from publicly available court documents and will undergo further validation to ensure its security. When releasing the dataset, we will provide usage limitations, appropriate scenarios, and guidelines. To minimize the risk of misuse, we will also implement access restrictions, including requiring applicants to provide their real identities.

### v2

@@ Abstract
Legal claims refer to the plaintiff’s demands in a case and are essential to guiding judicial reasoning and case resolution. While many works have focused on improving the efficiency of legal professionals, the research on helping non-professionals (e.g., plaintiffs) remains unexplored. This paper explores the problem of legal claim generation based on the given case’s facts. First, we construct ClaimGen-CN, the first dataset for Chinese legal claim generation task, from various real-world legal disputes. Additionally, we design an evaluation metric tailored for assessing the generated claims, which encompasses two essential dimensions: factuality and clarity. Building on this, we conduct a comprehensive zero-shot evaluation of state-of-the-art general and legal-domain large language models. Our findings highlight the limitations of the current models in factual precision and expressive clarity, pointing to the need for more targeted development in this domain. ClaimGen-CN dataset is available at: https://github.com/JosieZhou00/ClaimGen-CN.
@@ 4.1.2 Structured Data
Causes of Action Number
Private Lending Disputes 37,781
Motor Vehicle Traffic Accident Liability Disputes 24,025
Sales Contract Disputes 17,487
Financial Loan Contract Disputes 13,231
Contract Disputes (General) 8,906
Catering Service Contract Disputes 589
Trademark Infringement Disputes 66
Child Support Disputes in Cohabitation Relationships 70
Bank Card Disputes 80
Child Support Disputes 39
Table 3: Cause of action distribution.
Court Number
Shanghai Pudong New District People’s Court 3,452
Shanghai Minhang District People’s Court 2,329
Shanghai Baoshan District People’s Court 1,620
Shanghai Jing’an District People’s Court 1,470
Shanghai Qingpu District People’s Court 1,390
Ningbo Maritime Court 1
Yueyang Intermediate People’s Court, Hunan Province 1
Nantong Intermediate People’s Court, Jiangsu Province 1
Lixian People’s Court 1
Ankang Intermediate People’s Court, Shaanxi Province 1
Table 4: Court distribution.
Year Number
2010 7
2012 1
2013 3
2014 7
2015 21
2016 12,720
2017 47,749
2018 47,885
2019 38,966
2020 26,160
2021 28,656
2022 5,573
Table 5: Case start year distribution.
Statistic Fact Length Request Length Verdict Length
Mean 353.4 134.3 605.8
Std Dev 226.1 84.3 463.8
Min 7.0 14.0 103.0
Max 7,350.0 5,724.0 36,593.0
Table 6: Text length distribution.
@@ 4.2 Dataset Characteristics
Comprehensive. Beyond the general comparison, we further provide detailed distributions of ClaimGen-CN to demonstrate its comprehensiveness. As shown in Table 3, private lending disputes constitute the largest share, followed by motor vehicle traffic accident liability disputes and sales contract disputes, while a number of relatively rare categories such as trademark infringement and child support disputes also appear. Table 4 illustrates the court distribution, where cases are concentrated in Shanghai courts such as Pudong New District and Minhang District People’s Courts, though a small number of judgments come from courts in other provinces. Regarding temporal coverage, Table 5 shows that the dataset spans from 2010 to 2022, with a sharp increase starting in 2016 and peaking around 2017–2018 before declining in recent years. Finally, Table 6 reports the distribution of fact, request, and verdict text lengths, reflecting significant heterogeneity across different sections of judgments.
@@ 6.2 Future Work in Legal Claim Generation
Although our dataset is constructed based on the Chinese legal system, the overall data construction workflow and the evaluation design are not specific to China. With appropriate adjustments, especially considering the role of precedents in common law countries, our approach can be applied to other legal systems as well.
@@ 7 Conclusion
The corpus we use is released by the Chinese government and has been anonymized wherever necessary [^4] Provisions of the Supreme People’s Court on Publication of Judgment Documents by the People’s Courts on the Internet 2016. http://gongbao.court.gov.cn/Details/415f49dd8baaa04b479d57af9616ef.html. Our dataset thus does not involve any personally private information. Besides, the corpus is in public domain and licensed for use within the legal scope [^5] See Article 3.3 of User Protocol of China Judgments Online. https://g.alicdn.com/onlineCourt/static/0.6.97/akan-wenshu-protocol.html. Specifically, sensitive personal identifiers (such as names, identification numbers, and contact details) were removed or masked by the data providers before public release. Our research strictly used these desensitized datasets and did not involve any additional processing that could lead to re-identification. Beyond relying on officially anonymized and desensitized documents, we emphasize that our research activities did not introduce or generate any new sensitive information.
@@
Measures for releasing the dataset: The dataset was solely used for academic purposes such as model evaluation and error analysis. No attempts were made to re-identify individuals, and no system or output of this work has been, or will be, deployed in real-world legal practice. We emphasize that AI systems in legal contexts must be subject to strict human oversight, and our work is intended to support research and understanding rather than to replace human judgment. Any potential application of similar technologies should therefore incorporate comprehensive oversight mechanisms to prevent misuse.
@@
Acknowledgements
We would like to thank the anonymous reviewers and the Area Chair for their constructive feedback and insightful suggestions. We are also sincerely thankful to Chen Yanwen for suggesting the name of the dataset and for providing valuable comments and revisions.